OpenStudy (anonymous):

Composite functions. Make up any two functions f(x) and g(x), defined in terms of the variable x. For example, you might define f(x) = x – 15 and g(x) = 3x2. (Use different examples.) Then provide the compositions of the two functions f(g(x)) and g(f(x)). Which composite function has the greatest value for x = 10? Your answer completely depends on the functions you select.

OpenStudy (bahrom7893):

ok make up two functions.

OpenStudy (bahrom7893):

Whatever you pick will work.

OpenStudy (anonymous):

like this? f(x)=x-3 g(x)=(5x-3)

OpenStudy (bahrom7893):

Yea, so f(g(x)) = g(x) - 3

OpenStudy (bahrom7893):

f(g(x)) = (5x-3) - 3 = 5x - 6

OpenStudy (bahrom7893):

g(f(x)) = (5*f(x) - 3) = (5*(x-3)-3)

OpenStudy (bahrom7893):

g(f(x)) = 5x-18 Plug in 10 instead of x into both equations and see what's larger.

OpenStudy (anonymous):

Thanks, Bahroom. You're the best!

OpenStudy (bahrom7893):

ty